# VISinger2+: End-to-End Singing Voice Synthesis Augmented by Self-Supervised Learning Representation

### v1

@@ Abstract
keywords:
Singing voice synthesis, self-supervised learning
@@ 1 Introduction
This study introduces an innovative framework, namely VISinger2+, that leverages the strengths of pre-trained SSL models (e.g., HuBERT [15] and MERT [16]) to enrich the spectral inputs used in VISinger2 [8]. We conduct comprehensive experiments with our methods in both single-singer and multi-singer scenarios, as well as in different languages, such as Japanese and Mandarin. Our findings show that the proposed VISinger2+ can produce better quality singing voices than the baseline VISinger2 in most subjective and objective metrics.
@@ 2.1 Overview of VISinger2 Framework
Figure 1: The proposed VISinger2+ architecture. Model details are discussed in Section 2.
@@ 3.1 Experimental Datasets
For these three corpora, we followed the recipe from Muskits-ESPnet [35] to split the data into training, validation, and test sets.
@@ 3.3 Experimental Setups
Figure 2: t-SNE Visualization of singer embeddings on synthesized audios from 4 singers in the ACE-Opencpop dataset.
@@ 4 Results
Model MCD $\downarrow$ F0 RMSE $\downarrow$ ST Acc $\uparrow$
VISinger2 5.735 0.088 65.94%
VISinger2+ (HuBERT) 5.687 0.092 66.40%
VISinger2+ (MERT) 5.750 0.088 66.15%
@@
Model MCD $\downarrow$ F0 RMSE $\downarrow$ ST Acc $\uparrow$ SECS $\uparrow$
VISinger2 5.399 0.135 63.62% 0.795
VISinger2+ (HuBERT) 5.234 0.140 65.36% 0.833
VISinger2+ (CN-HuBERT) 5.259 0.145 65.02% 0.836
VISinger2+ (MERT) 5.181 0.141 64.93% 0.829

### v2

@@ Abstract
Index Terms—  Singing voice synthesis, self-supervised learning
@@ 1 Introduction
This study introduces a new framework, namely VISinger2+, that leverages the strengths of pre-trained SSL models (e.g., HuBERT [15] and MERT [16]) to enrich the spectral inputs used in VISinger2 [8]. We conduct comprehensive experiments with our methods in both single-singer and multi-singer scenarios, as well as in different languages, such as Japanese and Mandarin. Our findings show that the proposed VISinger2+ can produce better quality singing voices than the baseline VISinger2 in most subjective and objective metrics. Our open-source implementation can be found at  [^1] https://espnet.github.io/espnet/recipe/svs1.html#visinger-2-plus-training.
@@ 2.1 Overview of VISinger2 Framework
Fig. 1: The proposed VISinger2+ architecture. Model details are discussed in Section 2.
@@ 3.1 Experimental Datasets
For these three corpora, we followed the recipe from Muskits-ESPnet [35] [36] to split the data into training, validation, and test sets.
@@ 3.4 Evaluation
    We use SingMOS [38] as an additional MOS evaluation for Ofuton-P and ACE-Opencpop datasets, which lack human MOS evaluations. SingMOS leverages a pre-trained MOS prediction model to generate MOS-annotated results for singing voices. This approach enables efficient and reliable subjective ratings for these datasets.
@@ 4 Results
Model MCD $\downarrow$ F0 RMSE $\downarrow$ ST Acc $\uparrow$ SingMOS $\uparrow$
VISinger2 5.735 0.088 65.94% 3.42 ($\pm$ 0.08)
VISinger2+ (HuBERT) 5.687 0.092 66.40% 3.48 ($\pm$ 0.08)
VISinger2+ (MERT) 5.750 0.088 66.15% 3.49 ($\pm$ 0.09)
G.T. - - - 3.53 ($\pm$ 0.10)
@@
[…]
VISinger2 5.399 0.135 63.62% 0.795 3.74 ($\pm$ 0.01)
VISinger2+ (HuBERT) 5.234 0.140 65.36% 0.833 3.82 ($\pm$ 0.01)
VISinger2+ (CN-HuBERT) 5.259 0.145 65.02% 0.836 3.78 ($\pm$ 0.01)
VISinger2+ (MERT) 5.181 0.141 64.93% 0.829 3.85 ($\pm$ 0.01)
G.T. - - - - 3.85 ($\pm$ 0.01)
Fig. 2: t-SNE Visualization of singer embeddings on synthesized audios from 4 singers in the ACE-Opencpop dataset.
@@
The SingMOS scores provide further insights into the perceptual quality of the synthesized singing voices. As seen in Table 2, the VISinger2+ (MERT) model achieved the highest SingMOS score of 3.49 on the Ofuton-P dataset, closely matching the ground truth score of 3.53. Similarly, Table 3 shows that the VISinger2+ (MERT) model also obtained the highest SingMOS score of 3.85 on the ACE-Opencpop dataset, equaling the ground truth score. These results indicate that the VISinger2+ models not only enhance objective metrics but also significantly improve the perceived quality of the synthesized singing voices.